A skin heals according to N(t)+N(o)e^-0.02t, where N is the number of square centimeters of unhealed skin t days after an injury, and N(o) is the number of square centimeter covered by the original wound. How many days (to the nearest tenth ofa day) will it take for 50% of the wound to heal?
Answer by stanbon(75887)

N(t)=N(o)e^-0.02t How many days (to the nearest tenth ofa day) will it take for 50% of the wound to heal?
50% of N(o) = 0.5N(o)
0.5N(o) = N(o)e^(-0.02t)
0.5 = e^(-0.2t)
Take the natural log to get:
-0.2t=ln(0.5)
-0.2t=-0.6931...
t=3.5 days
Cheers,
Stan H.